### v1

@@ 3.1.1.3   Synthetic Data and Model Training
The DGM can also be used to generate various types of synthetic traffic data for traffic modeling and training purposes. These synthetic datasets enable researchers to simulate traffic scenarios and train models without relying solely on real-world data, which can be expensive and difficult to collect. In related research, instead of using basic models, most researchers combine various deep learning architectures or incorporate additional information to further improve performance (Chen et al., 2019; Ozturk et al., 2020; Wu et al., 2020; Chen et al., 2021c; Zhou et al., 2023; Rong et al., 2023b; Kumar et al., 2023; Devadhas Sujakumari and Dassan, 2023). For traffic data generation, Li et al. (2022a) proposed the Attentive Dual-Head Spatial-Temporal GAN (ADST-GAN) to generate crowd flows. The model integrates attentive temporal and spatial mechanisms by combining ConvLSTM, self-attention, and a dual-head discriminator within the GAN framework. These components capture complex spatial-temporal dependencies in crowd flow data and mitigate overfitting, ensuring that the synthetic data generated closely mimics real-world traffic patterns. To address the complexity of modeling Origin-Destination (OD) networks, Rong et al. (2023a) introduced a large-scale OD generation method using a graph-denoising diffusion model. This approach simplifies the generation process by decomposing it into two stages: first, generating the network topology, and second, generating the edge weights. By tackling these stages separately, the model accurately generates realistic OD data, making the analysis of large-scale OD networks more manageable. For model training, Dewi et al. (2022) applied a Deep Convolutional GAN (DCGAN) to generate synthetic traffic sign images. This approach captures the distribution of real traffic sign images and produces new samples to augment the original dataset. The researcher further combined these synthetic images with real ones to improve the performance of deep recognition models, such as CNN and ResNet 50. Additionally, Zhu et al. (2022) employed the Spatio-Temporal Dependencies GAN (STDGAN) to generate high-quality synthetic traffic volume data for prediction tasks. Specifically, STDGAN captures implicit variation patterns in traffic volume based on information from traffic speed and occupancy data. Experiment results indicate that enriching the training dataset with the synthetic volume data led to more robust and accurate performance in prediction models. These examples demonstrate the ability of DGMs to generate various traffic data that support traffic management and enhance model training, offering valuable resources for transportation research and applications.
@@ 3.1.1.4   Anomaly Data Generation
Anomaly data generation is crucial for creating rare or unexpected traffic situations, which are essential for testing and evaluating systems under extreme conditions. Generating such data helps develop robust models that can handle these rare events effectively. In this research field, GAN-based models are widely adopted due to their flexible and high performance in generating synthetic data for anomaly situations (Islam et al., 2021; Zarei and Hellinga, 2021; Li et al., 2024a; Chen et al., 2024). For instance, Cai et al. (2020) utilized a DCGAN to generate synthetic crash data from learning the distribution of crash-related traffic data. The synthetic data was then combined with real data to create a balanced dataset, which was used to train various crash prediction algorithms, including Logistic Regression, SVM, ANN, and CNN. Similarly, Chen et al. (2022d) employed two GAN variants—Conditional GAN (CGAN) and Gaussian Discriminant Analysis GAN (GDAGAN)—to tackle the challenge of imbalanced traffic collision datasets. These generative models generated additional samples for underrepresented classes, thereby balancing the dataset and enhancing the performance of classifiers trained on them. Huo et al. (2021) introduced the Text-to-Traffic GAN (T2GAN), which generates realistic traffic situations by integrating traffic data with semantic information from social media. In the proposed model, CGAN was used to generate traffic scenarios based on text inputs from social media, which enables the creation of varied and contextually rich traffic situations. To tackle data imbalance in traffic crash severity modeling, Chen et al. (2024) developed a method based on Conditional Tabular GAN (CTGAN) combined with Random Under-sampling (CTGAN-RU). This approach generated synthetic crash data while accounting for both continuous and discrete characteristics of the imbalanced dataset. The integration of various conditions helped create a more balanced training set and improving the accuracy and reliability of severity prediction models. These studies demonstrate the importance of using DGMs to generate anomaly data, which helps address the imbalanced data problem and improves the robustness of traffic management and safety systems.
@@ 3.1.2   Trajectory Generation
Trajectory generation creates realistic movement patterns for vehicles and pedestrians, which is essential for developing accurate traffic simulations and urban planning tools. By leveraging DGMs, researchers can produce synthetic trajectories that mimic real-world behaviors of traffic agents, enabling more robust analysis and testing of transportation systems. These generated trajectories help understand traffic dynamics, evaluate transportation policies, and enhance infrastructure design. These trajectories can be further classified based on their scale and scope, such as micro and macro scales. Micro-scale trajectories focus on detailed, short-term movement patterns of individual agents in localized areas, like intersections, while macro-scale trajectories represent broader, long-term traffic flow across extensive road networks or regions. The classification into micro and macro scales significantly impacts their use cases, such as collision avoidance or urban planning, and determines the complexity of the problems they address.
@@ 3.1.2.1   Micro-scale Trajectory Generation
Micro-scale trajectory generation focuses on producing detailed, short-term predictions for localized areas, such as intersections or specific road segments. These trajectories are critical for applications requiring precise and immediate responses, like collision avoidance, maneuver planning in dense urban settings, and real-time navigation adjustments. Numerous studies have leveraged various DGMS to produce realistic micro-scale trajectories (Krajewski et al., 2018; Ding et al., 2019; Krajewski et al., 2019; Gong et al., 2023; Yang et al., 2024). For example, Agarwal et al. (2020) used normalizing flows in generating trajectories for Autonomous Vehicles (AVs) through the FlowPlan method. This approach employs a Conditional VAE (CVAE) to encode human expert trajectories into a low-dimensional latent space. A Neural Autoregressive flow (NAF) then maps this latent space to a Boltzmann distribution over the AV control space, effectively capturing complex, multimodal distributions of feasible control trajectories conditioned on the scene context. Rempe et al. (2023) used a guided diffusion model to generate realistic pedestrian trajectories. The proposed model integrates with a physics-based humanoid controller, forming a closed-loop system capable of simulating large crowd movements across various scenarios. By combining the guided diffusion model with physical simulation, it accurately generates pedestrian trajectories that reflect real-world behavior. Similarly, Niedoba et al. (2024) employed a diffusion-based generative model to simulate realistic traffic scenarios involving multiple interacting agents. To be more specific, it diffuses the trajectories of all agents in a scene jointly, based on past, present, or future state observations. This method allows for the creation of diverse and realistic traffic scenarios, including rare, safety-critical events. These studies highlight how DGMs can produce highly detailed and realistic micro-scale trajectories, which are essential for applications that require precise and immediate response capabilities in traffic systems.
@@ 3.1.2.2   Macro-scale Trajectory Generation
Macro-scale trajectory generation covers broad spatial and temporal scopes, including extensive areas such as entire road networks or large regions over longer periods. The trajectories are crucial for strategic planning, traffic management, and long-term behavioral analysis. By examining origin-destination pairs, travel time distributions, and overall traffic flow patterns, macro-scale trajectories provide valuable insights into larger trends and systemic issues. The information is vital for urban planning, optimizing public transportation systems, and enhancing the efficiency of traffic networks. Numerous studies have focused on generating macro-scale trajectories with conditional information based on the enhanced DGM by various deep-learning models. (Chen et al., 2021c, b; Zhang et al., 2022b; Zhu et al., 2024a; Wei et al., 2024; Wang and Kankanhalli, 2024; Zhu et al., 2024b). For example, Xiong et al. (2023) introduced TrajSGAN, a model designed to generate human mobility trajectories for urban behavior analysis and decision-making. It integrates an attention-based generator for trajectory location prediction with a rollout module and a CNN-based discriminator to assess the spatial structure of the generated trajectories. This integration ensures that the results are realistic and semantically guided. TrajSGAN effectively reduces divergence in spatial-related metrics and can be utilized to study epidemic diffusion with high accuracy. Similarly, Choi et al. (2021a) developed TrajGAIL, a GAIL framework for generating urban vehicle trajectories. The model combines the capabilities of GAIL with RNNs to learn the underlying distributions of urban vehicle trajectory data. TrajGAIL allows for the generation of synthetic vehicle trajectories that closely resemble real-world patterns, even from limited observations. Sun and Kim (2023) introduced MAGAIL-VL, a data-driven, network-wide traffic simulation framework considering both Vehicle and Link agents. Specifically, MAGAIL-VL learns the distribution of vehicle movements and link states from observed data, enabling the generation of realistic traffic scenarios across an entire network. Recently, diffusion-based models have become popular for generating macro-scale trajectories (Zhu et al., 2024a; Wei et al., 2024; Wang and Kankanhalli, 2024; Zhu et al., 2024b). For example, Zhu et al. (2024a) introduced a novel approach DiffTraj, which generates high-quality synthetic GPS trajectories using a spatial-temporal diffusion probabilistic model. DiffTraj model combines residual blocks and multi-scale feature fusion in U-Net architecture to accurately predict noise levels during the reverse denoising process, enhancing its performance in capturing complex spatial-temporal dependencies and generating realistic synthetic trajectories. Similarly, Wei et al. (2024) presented Diff-RNTraj, a diffusion-based model designed to generate trajectories that are both geographically accurate and adhere to the logical constraints of the road network. The model combines a continuous diffusion framework with a pre-training strategy to manage the hybrid nature of road network-constrained trajectory data, resulting in realistic and usable synthetic trajectories. These studies highlight the diverse applications of DGMs in producing macro-scale trajectories. They provide essential tools and insights for improving transportation systems through strategic planning and management.

### v3

@@ 3.1.1.3   Synthetic Data and Model Training
The DGM can also be used to generate various types of synthetic traffic data for traffic modeling and training purposes. These synthetic datasets enable researchers to simulate traffic scenarios and train models without relying solely on real-world data, which can be expensive and difficult to collect. In related research, instead of using basic models, most researchers combine various deep learning architectures or incorporate additional information to further improve performance (Chen et al., 2019; Zhang et al., 2020a; Ozturk et al., 2020; Wu et al., 2020; Chen et al., 2021c; Zhou et al., 2023; Rong et al., 2023b; Kumar et al., 2023; Devadhas Sujakumari and Dassan, 2023). For example, Rong et al. (2023a) introduced a large-scale OD generation method using a graph-denoising diffusion model. This approach simplifies the generation process by decomposing it into two stages: first, generating the network topology, and second, generating the edge weights. By tackling these stages separately, the model accurately generates realistic large-scale OD data. In a similar way, Rong et al. (2023b) introduced an Origin-Destination Generation Networks (ODGN), a physics-informed machine learning framework for generating OD by combining Multi-view Graph Attention Networks (MGAT) to extract urban features and a gravity-guided predictor to estimate mobility flows between regions. The model employs a cGAN training strategy with a random walk sampling discriminator to capture overall network topology, producing OD networks that closely match real-world properties. In addition, Zhou et al. (2023) introduced KSTDiff, a Knowledge-enhanced Spatio-temporal Diffusion model that generates dynamic urban flow data for regions lacking historical records by leveraging an Urban Knowledge Graph (UKG) and a region-customized diffusion process guided by a learnable volume estimator. The framework can also be adapted for urban flow prediction, achieving competitive performance compared to state-of-the-art methods.
For model training, Ozturk et al. (2020) introduced a GAN-based traffic simulator that generates realistic and stochastic vehicle trajectories from real data, which are then used to train reinforcement learning agents. These generated data help RL agents train in a more realistic environment and significantly improve their generalization capabilities compared to agents trained on simple rule-based simulators. Dewi et al. (2022) applied a Deep Convolutional GAN (DCGAN) to generate synthetic traffic sign images. This approach captures the distribution of real traffic sign images and produces new samples to augment the original dataset. The researcher further combined these synthetic images with real ones to improve the performance of deep recognition models, such as CNN and ResNet 50. Additionally, Zhu et al. (2022) employed the Spatio-Temporal Dependencies GAN (STDGAN) to generate high-quality synthetic traffic volume data for prediction tasks. Specifically, STDGAN captures implicit variation patterns in traffic volume based on information from traffic speed and occupancy data. Experiment results indicate that enriching the training dataset with the synthetic volume data led to more robust and accurate performance in prediction models.
The provided examples demonstrate the ability of DGMs to generate various traffic data that support traffic management and enhance model training, offering valuable resources for transportation research and applications. One key limitation of using DGMs is ensuring that the synthetic data accurately reflects the complex, dynamic, real-world environment. In general, the generated data fails to fully capture the intricate patterns, dynamic shifts, and non-stationary behaviors seen in real scenarios, which can introduce biases into the model training process. This limitation not only affects the quality of the synthetic data but also complicates its integration with real data. Determining the optimal balance is another critical issue in model training. Specifically, an excess of synthetic data may overwhelm the true data distribution, while too little might not adequately address data scarcity. In summary, these challenges can hinder the model’s ability to generalize and perform effectively under dynamic conditions.
@@ 3.1.1.4   Anomaly Data Generation
Anomaly data generation is crucial for creating rare or unexpected traffic situations, which are essential for testing and evaluating systems under extreme conditions. Generating such data helps develop robust models that can effectively handle imbalanced datasets. In this research field, DGMs have been highly adapted to augment imbalanced traffic data. This augmentation not only addresses the scarcity of anomalous events in real-world datasets but also plays a pivotal role in enhancing the accuracy and resilience of prediction and classification models used in traffic safety and management applications.(Islam et al., 2021; Zarei and Hellinga, 2021; Li et al., 2024a; Chen et al., 2024). For instance, Cai et al. (2020) utilized a DCGAN to generate synthetic crash data from learning the distribution of crash-related traffic data. The synthetic data was then combined with real data to create a balanced dataset, which was used to train various crash prediction algorithms, including Logistic Regression, SVM, ANN, and CNN. Similarly, Chen et al. (2022d) compared several data augmentation methods including Synthetic Minority Oversampling Technique (SMOTE), GAN, conditional GAN (cGAN), and Gaussian Discriminant Analysis GAN (GDAGAN) to address the imbalance in traffic collision datasets. These models generated additional samples for underrepresented classes, thereby balancing the traffic collision dataset, and enhancing the performance of classifiers trained on them. Furthermore, Chen et al. (2024) developed a method that integrates Conditional Tabular GAN with Random Under-sampling (CTGAN-RU) to generate synthetic crash data. This approach accounts for both the continuous and discrete characteristics of imbalanced crash datasets by incorporating various conditions, thereby creating a more balanced training set and improving the accuracy and reliability of crash severity prediction models. Huo et al. (2021) introduced the Text-to-Traffic Generative Adversarial Network (T2GAN), a novel framework that generates realistic traffic situations by fusing traditional traffic data with semantic information extracted from social media. In T2GAN, a pre-trained GloVe model encodes the traffic-related text into semantic features that are integrated into the GAN training process, with a global-local loss employed to align the modalities and produce varied, contextually accurate traffic scenarios. The current studies underscore the importance of using DGMs for generating anomaly data to address imbalanced datasets and enhance the robustness of traffic management and safety systems. However, current models face significant challenges. In particular, they struggle to effectively manage imbalanced crash data that often includes missing values, heteroscedasticity, noise, and small sample sizes—characteristics common in real-world datasets. This highlights the need for further research to refine these methods and ensure that synthetic data generation more accurately reflects the complexities of actual traffic anomalies.
@@ 3.1.2   Trajectory Generation
Trajectory generation creates realistic movement patterns for vehicles and pedestrians, which is essential for developing accurate traffic simulations and urban planning tools. By leveraging DGMs, researchers can produce synthetic trajectories that mimic real-world behaviors of traffic agents, enabling more robust analysis and testing of transportation systems. These generated trajectories help understand traffic dynamics, evaluate transportation policies, and enhance infrastructure design. These trajectories can be further classified based on their scale and scope, such as micro and macro scales. Micro-scale trajectories focus on detailed, short-term movement patterns of individual agents in localized areas, like intersections, while macro-scale trajectories represent broader, long-term traffic flow across extensive road networks or regions. The classification into micro and macro scales significantly impacts their use cases, such as collision avoidance or urban planning, and determines the complexity of the problems they address. Table 3 presents an overview of recent studies in trajectory generation, organized by generation scale, task, model employed, dataset, and evaluation metrics.
Table 3: A summary of recent studies in trajectory generation using DGM
Generation Scale Author Task Model Dataset Evaluation Metrics
Micro Krajewski et al. (2018) GPS-level Generation (High Sampling) TrajGAN, TrajVAE HighD MSE
Krajewski et al. (2019) GPS-level Generation (High Sampling) BézierVAE HighD RMSE, Normalized Variance-of-Differences
Ding et al. (2019) GPS-level Generation (High Sampling) MTG ($\beta$-VAE-based) UMTRI Driving Encounter Dataset Rule-specific Metrics, Disentanglement Metric
Bhattacharyya et al. (2022) GPS-level Generation (High Sampling) GAIL NGSIM RMSE, Rule-specific Metrics
Chen et al. (2022a) GPS-level Generation (High Sampling) GAIL HighD RMSE
Ma et al. (2023) GPS-level Generation (High Sampling) PICGAN NGSIM MSE, Rule-specific Metrics
Yu et al. (2024) GPS-level Generation (High Sampling) TDS-GAN Closed Test Site from Highway Research Institute of the Ministry of Transport of China RMSE, MAPE, MAE, Rule-specific Metrics
Macro Choi et al. (2021a) Link-level Generation TrajGAIL Aimsun Simulation, Seoul Taxi DTG BLEU, METEOR, JSD
Sun and Kim (2023) Link-level Generation MAGAIL-VL pNEUMA BLEU, JSD, MAPE, NMAE
Xiong et al. (2023) Grid-level Generation TrajSGAN MTL Trajet Dataset JSD-based Metrics, Recall, $R^{2}$
Zhu et al. (2024a) GPS-level Generation (Low Sampling) DiffTraj Chengdu & Xi’an GPS Trajectory Dataset Density Error (JSD-based), Trip Error (JSD-based), Length Error, Pattern Score (F1-Score-based), MAE, MSE, RMSE
Wei et al. (2024) Link-level Generation Diff-RNTraj Porto & Chengdu Taxi Datasets JSD, Road Segment Connectivity
@@ 3.1.2.1   Micro-scale Trajectory Generation
Micro-scale trajectory generation involves producing detailed agent movements, closely focusing on behaviors within localized areas like intersections or specific road segments. It focuses on generating detailed vehicle dynamics, such as acceleration, deceleration, and lane-changing. The DGMs are particularly effective for this task as they efficiently learn complex and varied behaviors directly from real-world data, producing realistic trajectories. These accurate and detailed trajectories significantly enhance applications such as collision avoidance, precise maneuver planning, and road safety (Kuefler et al., 2017; Krajewski et al., 2018; Ding et al., 2019; Krajewski et al., 2019; Demetriou et al., 2020; Zhou et al., 2020b; Zhang et al., 2022d; Gong et al., 2023; Dong et al., 2023; Singh et al., 2023; Ma and Qu, 2023; Shi et al., 2024). For example, Krajewski et al. (2018) used two DGMS, Trajectory GAN(TraGAN) and VAE(TraVAE), which generate realistic lane change trajectories by learning intuitive, disentangled latent parameters from real-world driving data. These models can accurately reproduce observed maneuvers while also synthesizing new trajectories to fill gaps in simulation datasets. To further solve the smoothness problem, Krajewski et al. (2019) integrated a Bézier-curve output layer with additional loss terms into TraVAE, resulting in BézierVAE. This model generates smooth trajectories in both the position and speed domains, enhancing trajectory modeling for safety validation of highly automated vehicles. Ding et al. (2019) introduced a Multi-Vehicle Trajectory Generator (MTG) that integrates a VAE framework with bi-directional Gate Recurrent Units (GRUs) in the encoder and a multi-branch GRU decoder to simulate realistic vehicle-to-vehicle encounters. The authors also propose a novel disentanglement metric to assess the model’s stability and interpretability. In another study, Bhattacharyya et al. (2022) employed GAIL with several extensions (PS-GAIL, RAIL, and Burn-InfoGAIL) to model and replicate human driving behavior in simulation. By addressing the challenges of stochasticity, multimodality, and latent variability in human driving behavior, the proposed methods generate realistic car-following and lane-changing behaviors, as validated on the NGSIM dataset. Chen et al. (2022a) developed a two-step hybrid driving model that combines model-based controllers with GAIL for traffic simulation. The model generates high-level driver traits used as parameters for low-level model-based controllers to simulate human-like driving in multi-agent traffic scenarios. Ma et al. (2023) developed the Physics-Informed Conditional GAN (PICGAN) to control Connected Autonomous Vehicles (CAVs) in mixed traffic environments. The proposed model integrates theoretical physics with a dual conditional GAN framework, which consists of an encoder-decoder generator and two discriminators. By training on both observed and simulated data, PICGAN demonstrates improved robustness and adaptability. Similarly, Yu et al. (2024) proposed a Theory-data Dual Driven Stochastically GAN (TDS-GAN) by integrating a physics-informed GAN with a two-dimensional Intelligent Driver Model (2D-IDM) to capture the stochastic and heterogeneous car-following behavior of human-driven vehicles in mixed traffic. In this framework, the GAN captures the inherent randomness and uncertainty of car-following, generating realistic, varied trajectories that enhance the predictive accuracy of the theory-driven 2D-IDM for short-term forecasts and macroscopic traffic flow simulations. DGMs have demonstrated an ability to produce highly detailed and realistic micro-scale trajectories, which are crucial for applications that demand precise and immediate responses in traffic systems. Despite these promising results, several limitations remain to be solved. First, fine-tuning the guidance parameters of DGMs is often sensitive and requires extensive calibration to balance between trajectory diversity and fidelity. Specifically, even minor deviations can significantly affect the model’s ability to replicate rare or extreme driving events, such as sudden braking or abrupt lane changes, which are critical for safety. Moreover, the performance of DGMs is highly dependent on the quality and representativeness of the training data and any deficiencies in data coverage can compromise the model’s accuracy and its generalizability to unseen scenarios.
@@ 3.1.2.2   Macro-scale Trajectory Generation
Macro-scale trajectory generation covers broad spatial and temporal scopes, including extensive areas such as entire road networks or large regions over longer periods. Such trajectories are essential for strategic planning, traffic management, and long-term behavioral analysis. By examining OD pairs, travel time distributions, and overall traffic flow patterns, macro-scale trajectories provide valuable insights into larger trends and systemic issues. The information is crucial for urban planning, optimizing public transportation systems, and enhancing the efficiency of traffic networks. Numerous studies have applied DGMs to generate such macro-scale trajectories, offering valuable insights for strategic decision-making in transportation engineering.(Chen et al., 2021c, b; Zhang et al., 2022b; Zhu et al., 2024a; Wei et al., 2024; Wang and Kankanhalli, 2024; Zhu et al., 2024b). For example, Choi et al. (2021a) developed TrajGAIL, a GAIL framework for generating link sequences of urban vehicle trajectories. The model combines the capabilities of GAIL with RNNs to learn the underlying distributions of urban vehicle trajectory data. TrajGAIL allows for the generation of synthetic vehicle trajectories that closely resemble real-world patterns, even from limited observations. Similarly, Sun and Kim (2023) introduced MAGAIL-VL, a data-driven, network-wide traffic simulation framework considering both Vehicle and Link agents. Specifically, MAGAIL-VL learns the distribution of vehicle movements and link states from observed data, enabling the generation of realistic traffic scenarios across an entire network. In another study, Xiong et al. (2023) introduced TrajSGAN, a semantic-guiding adversarial network for urban trajectory generation that synthesizes human mobility trajectories at a grid scale. It integrates an attention-based generator for trajectory location prediction with a rollout module and a CNN-based discriminator to assess the spatial structure of the generated trajectories. This integrated framework effectively reduces divergence in spatial metrics and has been successfully applied to epidemic diffusion studies with high accuracy. Recently, diffusion-based models have gained popularity for generating fine-level urban trajectories. Recently, diffusion-based models have become popular for generating fine-level urban trajectories(Zhu et al., 2024a; Wei et al., 2024; Wang and Kankanhalli, 2024; Zhu et al., 2024b). For example, Zhu et al. (2024a) introduced DiffTraj, a novel approach for urban-scale trajectory generation that produces high-quality synthetic GPS trajectories using a spatial-temporal diffusion probabilistic model. DiffTraj leverages a U-Net architecture enhanced with residual blocks and multi-scale feature fusion to accurately estimate noise levels during the reverse denoising process. This design effectively captures complex spatial-temporal dependencies, enabling the generation of realistic synthetic trajectories that are well-suited for urban mobility analysis and preserving the statistical properties of real-world data. Similarly, Wei et al. (2024) presented Diff-RNTraj, a diffusion-based model designed to generate trajectories that are geographically accurate and adhere to road network constraints. The model uses a continuous diffusion framework combined with a pre-training strategy to handle the hybrid nature of road network–constrained trajectory data, resulting in realistic and usable synthetic trajectories. Even though the current studies have demonstrated significant advances in macro-scale trajectory generation, several challenges remain to be addressed. First, integrating external conditions—such as dynamic traffic states, diverse regional behaviors, and factors like departure times and local events—into the generative process is still limited, which restricts the models’ ability to accurately reflect real-time traffic dynamics. Second, ensuring that generated trajectories are physically plausible and strictly adhere to road network connectivity is challenging, often resulting in unrealistic or disconnected routes. Third, capturing the inherent stochasticity and heterogeneity of human mobility remains a major hurdle. Finally, the absence of unified evaluation metrics further complicates objective assessment and comparison among models, making it difficult to benchmark progress in urban-level trajectory generation using DGMs.